OpenStudy (anonymous):

Hi people, what is the solution of q capacitor diff equation : dq/dt+1/RC * q = 0

OpenStudy (anonymous):

Solve this:\[ y' + c y = 0 \]

OpenStudy (anonymous):

Let \(c= 1/RC\)

OpenStudy (anonymous):

This is an easy diff eq.

OpenStudy (anonymous):

\[ q'=-cq \]So \[ \frac{q'}{q}=-c \]And we get: \[ \ln(q) = -ct+ K \]Put it all together: \[ q(t) = Ke^{(1/RC) t} \]

OpenStudy (anonymous):

Notice \[ q(0) = Ke^{0} \implies K = q(0) = q_0 \]So \[ q(t) = q_0 e^{(1/RC)t} \]

OpenStudy (anonymous):

@wio and what is q0 ?

OpenStudy (anonymous):

It is \(q\) when \(t=0\).

OpenStudy (anonymous):

i can't understand why it's CEe^-t/yau dvc/dt+1/RC*vc=0 the same equation but the solutions is : Ee^-t/tau

OpenStudy (anonymous):

I might have forgotten a negative sign.

OpenStudy (turingtest):

yes, you dropped a negative sign, wio, and I'm not sure exactly what you mean @AntarAzri , the solution to the equation you posted is\[Q(t)=Q_0e^{-t/\tau}=C\mathcal E e^{-t/\tau}=CV(t)\]as you can see you can get an equation for the voltage across the capacitor as a function of time by cancelling \(C\) from both sides, which gives your second equation\[V(t)=\mathcal E e^{-t/\tau}\]note that you could have canceled the \(C\) before solving the equation, and instead solved\[\frac{dV}{dt}+{V\over RC}=0;~~~V(0)=\mathcal E\]